OpenStudy (anonymous):

Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4.

OpenStudy (anonymous):

i am 100% lost lol

OpenStudy (anonymous):

y(1)=8*1 y(2)=8*2=16 y(3)=8*3=24 y(4)=8*4=32

OpenStudy (anonymous):

f(x) x 8 1 16 2 24 3 32 4

OpenStudy (kymber):

 :D